# Does Gaussian Splatting need SFM Initialization?

### v2

@@ 1 Introduction
In the years since their introduction, Neural Radiance Fields (NeRF) [17] have emerged as a leading technique for 3D reconstruction from images and novel view synthesis, capturing the interest of researchers and practitioners alike. Despite its success, the application of NeRF in practical scenarios has been limited by its high computational demands. Due in part to this limitation, 3D Gaussian Splatting [9] has gained traction as a more efficient alternative, as it achieves real-time inference speeds by leveraging hardware rasterization.
However, the adoption of Gaussian Splatting is not without challenges. A significant hurdle is its dependence on careful initialization from point clouds, such as those from Structure-from-Motion (SfM).
The reliance on slow and computationally expensive SFM implementations such as COLMAP [25, 26], and comparatively poor performance of random initialization is a critical limitation that impedes the application of Gaussian Splatting in domains where a full SFM solution would be too expensive. This includes situations like SLAM sequences where accurate camera poses are available alongside point clouds too sparse to be useful for initialization, such as ORB-SLAM [19], or autonomous vehicle applications where navigation and localization are performed with other sensors, e.g., aircraft.
@@
Through extensive experiments, we find that a combination of better random initialization and structure guidance from NeRF are sufficient to match or exceed the quality reached from a COLMAP initialization, even on very large-scale, challenging scenes. This result could have significant implications for applications of Gaussian Splatting which do not require an SFM solution to estimate camera poses, such as autonomous vehicles with fused inertial/satellite navigation, or for which SFM solutions are not reliable. We also find that the required training time for the NeRF teacher model can be very low, on the order of 30 seconds – much faster than large SFM reconstructions, and even the Gaussian Splatting training itself.
@@ 2 Related Work
While significant advances have been made in improving the efficiency of models based on volume rendering, it is still a challenge to reach truly interactive inference speeds with them, due largely to the Monte Carlo integration required for each ray. An alternative approach to differentiable rendering that avoids this is rasterization, in which individual primitives are projected into image space and written to a buffer in order of depth. One such method is Pulsar [12], which uses partially transparent spheres as a representation and demonstrates the expressive power of such primitive-based models. More recently, Gaussians were proposed as a primitive [9], and were shown to achieve both very high-quality image reconstruction and better than real-time inference speeds. This result sparked a significant number of follow-up works, ranging from modelling dynamic content [32, 30, 28, 11], to generative modelling [34, 2, 27]. There were also improvements made to the handling of primitive scale in rendering to avoid aliasing [29, 31].
@@
Other differentiable rendering works have also recognized the potential of depth regularization and supervision in improving novel view synthesis results. Works such as RegNeRF [21], LOLNeRF [23], and Mip-NeRF 360 [1] all proposed some form of regularization of the density sampled along a ray in order to improve the recovered 3D structure. Other works [20, 24, 5] have explored directly supervising depth to improve the final rendered results.
@@ 3 Method
However, unlike NeRF, the point-based nature of Gaussian Splatting models is less able to converge from a random initialization far from the final distribution of primitives. For this reason, the original Gaussian Splatting method employed discrete splitting, cloning, and pruning heuristics, which interrupt the continuous optimization to take larger steps toward an optimal configuration, as defined by tuned thresholds. This approach is at least partially effective in closing the gap between Gaussian Splatting and NeRF in terms of scene structure recovery, but for large-scale, complex scenes, the original publication [9] still reported significant gains when initializing from an SFM point cloud versus a uniform random distribution.
In the remainder of this section, we will describe the potential strategies for closing this gap which we chose to evaluate.
@@ 3.2 Initialization from a volumetric model
Beyond seeking an “improved” random initialization, we also look for ways that the superior structure recovery of volumetric NeRF models can be leveraged to bootstrap the training of Gaussian Splatting. We approach this by using a probabilistic interpretation of NeRF density in volume rendering which sees volume rendering weights as a distribution of “surface” locations, i.e. where a ray is likely to terminate [7]. This differential [^1] Note we use an unbounded differential value so as to define a continuous PDF, rather than integrating over a local interval as in [7]. “surfaceness” value at a depth $t$ along some ray $\mathbf{r}$ is defined as:
@@ 3.4 Hyper-parameters
The original Gaussian Splatting implementation depends on a number of parameters which are tuned based on the original photometric loss formulation. As such, it is necessary to take care in setting parameters when changing the loss formulation. We found through experimentation that including a depth term in the loss function alters the training dynamics of pruning, splitting and cloning, which can lead to an excessive number of Gaussians being instantiated if the wrong parameters are used. For all real data scenes, we use an exponential decay schedule for the depth loss weight $\lambda$ with an initial value of 0.9, and a decay rate of 0.9 every 100 iterations. For synthetic scenes, we set the initial $\lambda$ to 0.95 and the decay value to 0.8. We also observed better results by changing the start of densification to iteration 100 rather than 500.
@@ 4 Experiments
(a) Ground truth
(b) COLMAP
(c) NerfAcc
(d) NerfAcc + depth
(e) Random
(f) Random + depth
Figure 3: Qualitative results on a scene from the Mip-NeRF 360 dataset [15]. We observe that the denser initialization from NeRF, and the structure guidance both lead to decreased loss of detail and thin geometry. Please zoom in to see details.
(a) Ground truth
(b) COLMAP
(c) NerfAcc
(d) NerfAcc + depth
(e) Random
(f) Random + depth
Figure 4: Qualitative results on a scene from the OMMO dataset [15]. We find that the COLMAP and random initializations allocate fewer points to the surface of the water, which leads to over-smoothed reconstruction compared to the NeRF-based initialization and models with depth supervision. Please zoom in to see details.
(a) COLMAP Initialization
(b) Ours
(c) COLMAP Initialization
(d) Ours
(e) COLMAP Initialization
(f) Ours
(g) COLMAP Initialization
(h) Ours
Figure 5: Additional qualitative results from the OMMO dataset [15] and Mip-NeRF 360. Here we visualize only the result of training with COLMAP (Left) and our best performing model with NeRF initialization + depth supervision (Right), to highlight the total contribution of the strategies we evaluated. Please zoom in to see details.
(a) NerfAcc
(b) Gaussian Splatting
(c) NerfAcc
(d) Gaussian Splatting
(e) NerfAcc
(f) Gaussian Splatting
@@ 4.2 Analysis of random initialization
Table 3: Volumetric initialization (OMMO). We evaluate test PSNR on the OMMO dataset [15] when initializing the Gaussian Splatting model from INGP models trained for different amounts of time.
@@
As reported in Tables 2 and 3, we find that these initializations near-universally outperform purely random initialization, and for some scenes outperform COLMAP initialization, even when using NeRF models trained for only 5000 iterations (about 30 seconds on our RTX 4090 GPU).
@@ 4.4 Depth distillation from NeRF
Table 6: Structure distillation (NeRF synthetic). We evaluate test PSNR on the NeRF synthetic dataset [17] when applying our depth loss for structure distillation from INGP models trained for 5k iterations.
@@
Initialization Lego Ship Hotdog Materials Ficus Chair Drums Mic Average
Random (From [9]) 35.78 30.81 37.72 30.00 34.87 35.83 26.15 35.36 33.31
Random (Re-run) 36.38 31.80 38.14 30.51 35.50 35.73 26.30 36.75 33.89
Random + Depth Loss 36.45 31.74 38.29 30.43 35.04 35.43 26.10 36.80 33.79
Finally, we experiment with integrating direct depth guidance from the NeRF model into the Gaussian Splatting training pipeline (see Figure 6 for examples). Using the loss formulation in (7), we directly supervise predicted depth values from Gaussian Splatting with the pre-trained NeRF model, which enables more fine-grained transfer of structure. It also helps avoid erroneous pruning of geometry as shown in Figure 3.
We additionally check that our depth supervision strategy is not harmful in cases where standard Gaussian Splatting is already sufficient by testing it on the NeRF synthetic dataset [17]. This dataset is object-centric, with blank backgrounds, and performs well with only a random initialization to the bounding box of the object. As shown in Table 6, we find that there is minimal difference in average between the depth-supervised and standard training on this dataset, suggesting that inclusion of depth distillation is not harmful in cases where it is not strictly necessary.
@@ 5 Conclusion
We performed an analysis of several approaches for avoiding SFM initialization in training of Gaussian Splatting. Through our experiments, we found that random initialization can perform better than expected based on the results of the original Gaussian Splatting paper. We also found that use of a small amount of time spent pre-training a NeRF model can enable results on par with or better than those achieved by COLMAP initialization by providing both a dense initial point cloud and depth supervision throughout training.
@@
The most immediate limitation in applying the strategies we have described to avoid requiring SFM solutions for Gaussian Splatting is that for many use cases, SFM is still the most reliable way of obtaining the camera calibration necessary to train. While our results suggest that there can still be value to leveraging NeRF pre-training when SFM points are available, it would be ideal to combine with other pipeline components capable of replacing SFM camera estimates.
@@
Appendix A Appendix Section
@@ A.1 Results on additional data
COLMAP (Re-run) 29.85 / 0.8946 / 0.1585 28.96 / 0.8876 / 0.1774 29.41 / 0.8911 / 0.1680
NerfAcc @ 5k 30.33 / 0.8962 / 0.1530 29.41 / 0.8929 / 0.1707 29.87 / 0.8946 / 0.1618
$50^{3}$ Random + Depth Loss 29.92 / 0.8874 / 0.1628 29.10 / 0.8863 / 0.1838 29.51 / 0.8868 / 0.1733
NerfAcc @ 5k + Depth Loss 29.90 / 0.8883 / 0.1581 29.54 / 0.8929 / 0.1717 29.72 / 0.8906 / 0.1649
A.2 Additional metrics
We additionally report the perceptual quality metrics SSIM and LPIPS [33] for the Mip-NeRF 360 [1] and OMMO [15] datasets. These results can be found in Tables 8, 9, 10, and 11.
@@ A.2 Additional metrics
A.3 Depth rendering with Gaussians
$G(x)=\textrm{exp}\left(-\tfrac{1}{2}x^{T}\Sigma^{-1}x\right),$ (8)
By expanding the polynomial, we can rewrite this as:
$d=$ $-\frac{d^{T}\Sigma^{-1}o}{d^{T}\Sigma^{-1}d}.$ (12)
A.4 Scheduling of depth loss weight
We use the following schedule for the loss weight $\lambda$:
$\lambda(i)=\lambda_{init}\times d^{\frac{i}{S}},$ (13)

### v3

# Evaluating Alternatives to SFM Point Cloud Initialization for Gaussian Splatting

@@ 1 Introduction
In the years since their introduction, Neural Radiance Fields (NeRF) [22] have emerged as a leading technique for 3D reconstruction from images and novel view synthesis, capturing the interest of researchers and practitioners alike. Despite its success, the application of NeRF in practical scenarios has been limited by its high computational demands. Due in part to this limitation, 3D Gaussian Splatting [12] has gained traction as a more efficient alternative, as it achieves real-time inference speeds by leveraging rasterization.
[…]
@@
Through extensive experiments, we find that a combination of better random initialization and structure guidance from NeRF are sufficient to match or exceed the quality reached from a COLMAP initialization, even on very large-scale, challenging scenes. We apply this scheme to scenes with cameras estimated using a state-of-the-art SLAM method [2], and show that the results are competitive with those derived from COLMAP initialization while taking far less time overall to go from images to the final model.
@@ 2 Related Work
While significant advances have been made in improving the efficiency of models based on volume rendering, it is still a challenge to reach truly interactive inference speeds with them, due largely to the Monte Carlo integration required for each ray. An alternative approach to differentiable rendering that avoids this is rasterization, in which individual primitives are projected into image space and written to a buffer in order of depth. One such method is Pulsar [16], which uses partially transparent spheres as a representation and demonstrates the expressive power of such primitive-based models. More recently, Gaussians were proposed as a primitive [12], and were shown to achieve both very high-quality image reconstruction and better than real-time inference speeds. This result sparked a significant number of follow-up works, ranging from modelling dynamic content [41, 39, 36, 15], to generative modelling [44, 3, 35]. There were also improvements made to the handling of primitive scale in rendering to avoid aliasing [38, 40]. A concurrent work further views the training process of Gaussian Splatting as a sampling process [13], and attempts to remove initialization dependence, similarly to ours.
Gaussian Splatting SLAM
There have also been a number of methods proposed for using Gaussian Splatting as an internal representation for Simultaneous Localization And Mapping (SLAM) [37, 42], which enables solving camera estimation and the construction of a representation for novel view synthesis at the same time given a stream of RGB and depth data. One such recent work [20] also supports performing SLAM on RGB-only video – we compare the results of this method with the strategies we present for COLMAP-free training. Another similar work [9] supports RGB-only training, but has not yet made code available for comparison.
@@
Other differentiable rendering works have also recognized the potential of depth regularization and supervision in improving novel view synthesis results. Works such as RegNeRF [26], LOLNeRF [29], and Mip-NeRF 360 [1] all proposed some form of regularization of the density sampled along a ray in order to improve the recovered 3D structure. Other works [25, 30, 6] have explored directly supervising depth to improve the final rendered results. One concurrent work to ours [27] also explores initialization and supervision of Gaussian Splatting from NeRF models, with similar findings regarding the benefits of the improved initialization. Another concurrent work [7] also explores replacing COLMAP initialization, specifically with pre-trained camera pose and scene structure estimation networks.
@@ 3 Method
However, unlike NeRF, because Gaussians are bounded locally in practice despite their theoretically infinite support, Gaussian Splatting models are less likely to converge from a random initialization far from the final distribution of primitives. For this reason, the original Gaussian Splatting method employed discrete splitting, cloning, and pruning heuristics, which interrupt the continuous optimization to take larger steps toward an optimal configuration, as defined by tuned thresholds. This approach is at least partially effective in closing the gap between Gaussian Splatting and NeRF in terms of scene structure recovery, but for large-scale, complex scenes, the original publication [12] still reported significant gains when initializing from an SFM point cloud versus a uniform random distribution.
In the remainder of this section, we will describe the potential strategies for closing this gap which we empirically explore.
@@ 3.2 Initialization from a volumetric model
Beyond seeking an “improved” random initialization, we also look for ways that the superior structure recovery of volumetric NeRF models can be leveraged to bootstrap the training of Gaussian Splatting. We approach this by using a probabilistic interpretation of NeRF density in volume rendering which sees volume rendering weights as a distribution of “surface” locations, i.e. where a ray is likely to terminate [10]. This differential [^1] To define a continuous PDF, we use an unbounded differential value rather than integrating over a local interval as in [10]. “surfaceness” value at a depth $t$ along some ray $\mathbf{r}$ is defined as:
@@ 3.4 Hyper-parameters
The original Gaussian Splatting implementation depends on a number of parameters which are tuned based on the original photometric loss formulation. As such, it is necessary to take care in setting parameters when changing the loss formulation. We found through experimentation that including a depth term in the loss function alters the training dynamics of pruning, splitting and cloning, which can lead to an excessive number of Gaussians being instantiated if the wrong parameters are used. For all real data scenes, we use an exponential decay schedule for the depth loss weight $\lambda$ with an initial value of 0.9, and a decay rate of 0.9 every 100 iterations.
@@ 4 Experiments
Ground truth
COLMAP
NerfAcc
NerfAcc + depth
Random
Random + depth
Figure 3: Qualitative results on a scene from the OMMO dataset [19]. We find that the COLMAP and random initializations allocate fewer points to the surface of the water, which leads to over-smoothed reconstruction compared to the NeRF-based initialization and models with depth supervision. Please zoom in to see details. Please also refer to the interactive comparison in the supplementary material.
@@ 4.1 Datasets
For these initial experiments, we continue to use camera parameters estimated by COLMAP to avoid the confounding effect of different camera parameters on our analysis. To further demonstrate how our findings translate when SfM solutions are not available, we also test a fully COLMAP-free pipeline on two SLAM datasets: Replica [33] and TUM [34]. For these datasets, we apply the volumetric pre-training and guidance strategies using cameras estimated by ORB-SLAM3 [2], and compare the results to those using COLMAP cameras/initializations, as well as the results of a recent Gaussian Splatting SLAM method [20].
@@ 4.2 Analysis of random initialization
Table 3: Volumetric initialization (OMMO). We evaluate test PSNR on the OMMO dataset [19] when initializing the Gaussian Splatting model from INGP models trained for different lengths.
@@
As reported in Tables 2 and 3, we find that these initializations near-universally outperform purely random initialization, and for some scenes outperform COLMAP initialization, even when using NeRF models trained for only 5000 iterations (about 30 seconds on our RTX A6000 GPU).
@@ 4.4 Depth distillation from NeRF
Next, we experiment with integrating direct depth guidance from the NeRF model into the Gaussian Splatting training pipeline (see Figure 6 for examples). Using the loss formulation in (7), we directly supervise predicted depth values from Gaussian Splatting with the pre-trained NeRF model, which enables more fine-grained transfer of structure. It also helps avoid erroneous pruning of geometry as shown in Figure 5.
@@
4.5 COLMAP-free training
Table 6: SLAM camera estimates (TUM). We compare the performance of NerfAcc initialization and depth loss, using cameras estimated by ORB-SLAM3 [2], to that of standard Gaussian Splatting with COLMAP cameras and a recent Gaussian Splatting SLAM method on the TUM dataset [34]. Along with test PSNR, we report the total run-time of each pipeline from images to the trained model.
@@ 4.5 COLMAP-free training
Model Desk XYZ Long Average
COLMAP 27.51 | 0:15:58 30.04 | 7:38:42 28.19 | 1:14:06 28.58 | 3:02:55
MonoGS [20] 20.90 | 0:06:46 22.38 | 0:19:58 22.41 | 0:16:01 21.90 | 0:14:15
NerfAcc @ 5k 24.91 | 0:07:05 27.01 | 0:09:47 28.15 | 0:07:36 26.69 | 0:08:09
NerfAcc @ 5k + Depth Loss 25.01 | 0:08:22 27.24 | 0:18:30 28.27 | 0:12:10 26.84 | 0:13:01
Table 7: SLAM camera estimates (Replica). We compare the performance of NerfAcc initialization and depth loss, using cameras estimated by ORB-SLAM3 [2], to that of standard Gaussian Splatting with COLMAP cameras and a recent Gaussian Splatting SLAM method on the Replica dataset [33]. Along with test PSNR, we report the total run-time of each pipeline from images to the trained model.
Scene
Model Office0 Office1 Office2 Office3 Office4 Room0 Room1 Room2 Average
COLMAP 44.65 | 0:38:33 44.20 | 0:38:51 40.54 | 1:01:20 40.61 | 0:52:18 38.46 | 0:35:05 39.40 | 1:13:51 32.59 | 2:55:00 41.66 | 0:49:44 40.26 | 1:05:35
MonoGS [20] 34.42 | 0:22:11 35.14 | 0:21:51 28.11 | 0:24:30 29.45 | 0:24:14 27.49 | 0:23:38 28.07 | 0:26:19 26.58 | 0:24:10 28.11 | 0:24:13 29.67 | 0:23:53
NerfAcc @ 5k 42.94 | 0:09:47 32.75 | 0:07:47 35.70 | 0:13:52 35.87 | 0:11:52 35.09 | 0:10:50 34.86 | 0:18:00 36.54 | 0:14:52 38.33 | 0:12:56 36.51 | 0:12:29
NerfAcc @ 5k + Depth Loss 43.03 | 0:19:00 33.17 | 0:17:51 35.76 | 0:22:52 35.80 | 0:21:51 35.06 | 0:20:52 34.88 | 0:25:58 36.49 | 0:22:55 38.32 | 0:21:49 36.56 | 0:21:38
Finally, we combine the best-performing strategies from the previous experiments with cameras estimated by a monocular SLAM method (ORB-SLAM3 [2]) which is much faster than COLMAP. We also compare to a recent method that combines monocular SLAM and Gaussian Splatting (MonoGS) to estimate cameras during training [20]. As shown in Tables 6 and 7, we find that the approaches we tested result in higher PSNR on average than MonoGS, and run significantly faster than the COLMAP-based pipeline, albeit with a somewhat lower PSNR, likely due to less accurate pose estimation. We include video renders of sample sequences in the Supplementary Material for comparison. The run times for the ORB-SLAM3 experiments include the SLAM tracking times themselves, about 40ms per frame, as well as the NerfAcc training and rendering time, about 300ms per frame if including depth supervision, or about 45ms per frame if only sampling a point cloud.
To be as fair as possible, we run COLMAP for these scenes with sequential matching, which only works for video input but results in much faster camera estimation. Despite this, we found COLMAP to still be extremely slow (20$\times$ on average on TUM compared to the NerfAcc init pipeline and up to 45×45$\times$45 × slower) for long sequences. We also reduce the size of the NerfAcc-generated point cloud to 50,000 points to more closely match the SFM point cloud, which is much smaller for these indoor non-360 scenes. All timed experiments are run on the same machine with an RTX A6000 GPU.
@@ 5 Conclusion
We performed an analysis of several approaches for avoiding SFM initialization in training of Gaussian Splatting. Through our experiments, we found that random initialization can perform better than expected based on the results of the original Gaussian Splatting paper. We also found that use of a small amount of time spent pre-training a NeRF model can enable results on par with or better than those achieved by COLMAP initialization by providing both a dense initial point cloud and depth supervision throughout training. We also tested these strategies with SLAM-estimated cameras and found that they can result in significant time savings across the entire pipeline.
@@
The most immediate limitation in applying the strategies we have described to avoid requiring SFM solutions for Gaussian Splatting is that, as our SLAM experiments show, SFM is still the most reliable way of obtaining the camera estimates necessary to train for many use cases. As such, careful consideration of the target application is needed to determine whether alternatives to SFM are available which can provide sufficiently accurate camera poses.
@@
6 Acknowledgements
This work was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant [2023-05617], NSERC Collaborative Research and Development Grant, the SFU Visual Computing Research Chair, Google, Digital Research Alliance of Canada, and Advanced Research Computing at the University of British Columbia.
Appendix A Appendix
COLMAP
Ours
COLMAP - Error Map
Ours - Error Map
Figure 4: Additional qualitative results from the OMMO dataset [19] and Mip-NeRF 360 [1]. Here we visualize only the result of training with COLMAP and our best performing model with NeRF initialization + depth supervision, to highlight the total contribution of the strategies we evaluated. Please zoom in to see details.
Ground truth
COLMAP
NerfAcc
NerfAcc + depth
Random
Random + depth
Figure 5: Qualitative results on a scene from the Mip-NeRF 360 dataset [19]. We observe that the denser initialization from NeRF, and the structure guidance both lead to decreased loss of detail and thin geometry. Please zoom in to see details. Please also refer to the interactive comparison in the supplementary material.
NerfAcc
Gaussian Splatting
@@ A.1 Results on additional data
A.2 Additional metrics
We additionally report the perceptual quality metrics SSIM and LPIPS [43] for the Mip-NeRF 360 [1] and OMMO [19] datasets. These results can be found in Tables 9, 10, 11, and 12.
A.3 Depth rendering with Gaussians
$G(x)=\textrm{exp}\left(-\tfrac{1}{2}x^{T}\Sigma^{-1}x\right),$ (8)
By expanding the polynomial, we can rewrite this as:
$d=$ $-\frac{d^{T}\Sigma^{-1}o}{d^{T}\Sigma^{-1}d}.$ (12)
A.4 Scheduling of depth loss weight
We use the following schedule for the loss weight $\lambda$:
$\lambda(i)=\lambda_{init}\times d^{\frac{i}{S}},$ (13)
@@ A.4 Scheduling of depth loss weight
COLMAP (Re-run) 29.03 / 0.8900 / 0.1763 30.10 / 0.9000 / 0.1587 29.56 / 0.8950 / 0.1675
NerfAcc @ 5k 29.53 / 0.8952 / 0.1684 30.66 / 0.9027 / 0.1524 30.09 / 0.8990 / 0.1604
$50^{3}$ Random + Depth Loss 29.38 / 0.8904 / 0.1816 30.06 / 0.8937 / 0.1615 29.72 / 0.8921 / 0.1716
NerfAcc @ 5k + Depth Loss 29.65 / 0.8965 / 0.1691 30.33 / 0.8946 / 0.1569 29.99 / 0.8956 / 0.1630
@@
A.5 Broader impact
We do not propose any new methods or technologies that would pose any novel risks to society which require mitigation. However, our analysis focuses on the lowest-cost ways to train novel view synthesis models, and the results of this analysis could be relevant in determining the risks posed by such technologies in cases where low barrier to entry could increase the likelihood of misuse as part of synthetic or deceptive media.
A.6 Dataset licenses
We use the following datasets: